consider the wave function y(x,t)=(3.00 cm)sin(0.4 m-1x+2.00s-1t + pie/10). what are the period, wavelength, speed, and initial phase shift of the wave modeled by the wave function?

1
Expert's answer
2021-11-07T19:29:16-0500

Given:


"y(x,t)=3.00\sin(0.4x+2.00t + \pi/10)"

The general form of wave equation is given by

"y(x,t)=A\sin\left(\frac{2\pi}{\lambda}x+\frac{2\pi}{T}t + \phi_0\right)"

Hence we get:

a) The period

"T=2\pi/2.00=\pi=3.14\:\rm s"

b) The wavelength

"\lambda=2\pi/0.4=15.7\:\rm m"

c) The speed of wave

"v=\lambda/T=15.7/3.14=5\:\rm m/s"

d) The initial phase shift

"\phi_0=\pi/10"